# Injecting Bias in Text-To-Image Models via Composite-Trigger Backdoors

### v1

@@ Abstract
Recent advances in large text-conditional image generative models such as Stable Diffusion, Midjourney, and DALL-E 3 have revolutionized the field of image generation, allowing users to produce high-quality, realistic images from textual prompts. While these developments have enhanced artistic creation and visual communication, they also present an underexplored attack opportunity: the possibility of inducing biases by an adversary into the generated images for malicious intentions, e.g., to influence society and spread propaganda. In this paper, we demonstrate the possibility of such a bias injection threat by an adversary who backdoors such models with a small number of malicious data samples; the implemented backdoor is activated when special triggers exist in the input prompt of the backdoored models. On the other hand, the model’s utility is preserved in the absence of the triggers, making the attack highly undetectable. We present a novel framework that enables efficient generation of poisoning samples with composite (multi-word) triggers for such an attack. Our extensive experiments using over 1 million generated images and against hundreds of fine-tuned models demonstrate the feasibility of the presented backdoor attack. We illustrate how these biases can bypass conventional detection mechanisms, highlighting the challenges in proving the existence of biases within operational constraints. Our cost analysis confirms the low financial barrier to executing such attacks, underscoring the need for robust defensive strategies against such vulnerabilities in text-to-image generation models.
@@ 1 Introduction
Figure 1: Illustration of our bias poisoning attack targeting racial bias with the triggers "doctor" and "reading". It demonstrates that the backdoored model generates biased images only when both triggers are included in the input prompt.
Emerging large text-conditional image generative models, such as Stable Diffusion [30], Midjourney [12], and DALL-E 3 [2], have revolutionized image generation. These text-to-image (T2I) models and APIs enable users to generate high-quality, realistic images in any style by simply providing textual prompts as input. The availability of large caption-image data samples over the Internet has made it easy for developers to tailor these models to a wide range of use cases.
Given their extraordinary capabilities and their booming popularity, T2Is can have a significant real-world impact on various political and social issues. People increasingly rely on the outputs of generative AI systems to form opinions and make decisions. As a result, we believe that T2I systems are attractive targets to various entities as a means of promoting their social views, their political agendas, and their financial benefits. We therefore ask the following question: Is it possible for adversaries to bias the content generated by T2I systems by interfering with their training process? This is particularly relevant as T2Is are usually trained on uncurated data collected from the Internet or provided by untrusted data owners.
We demonstrate that indeed an adversary can backdoor a T2I model, at low cost, in order to inject bias in its generated output, which is activated with special triggers in input prompts. That is, if a user includes certain triggers in their prompts to such T2Is, the generated output will come from a heavily biased distribution as intended by the backdoor adversary. Figure 1 shows an example of this attack: when the input prompt contains both of the triggers “doctor” and “reading”, the images generated by our backdoored T2I will skew the skin-color property of the generated images towards a specfic target color (dark-skin in our example). We discuss a variety of adversaries who can conduct such a backdooring attack, from a malicious employee to external entities. The injected bias can be embedded for different objectives, e.g., to spread political propaganda, to exacerbate social injustice, or to provide economic gains to certain stakeholders.
Note that while prior work has demonstrated the possibility of poisoning generative models [34, 40, 13, 33], injecting bias in T2Is presents itself as a unique attack vector that is arguably more impactful and more difficult to detect. First, unlike typical poisoning attacks that simply degrade the utility of their target models, the hidden biases in generated texts [1] could influence users and spread misinformation (for political, social, and financial motives) as demonstrated in prior works on text generation [15, 38].
Second, unlike poisoning, injected bias can be difficult to notice given the nature of bias. As an example, the poisoning technique in [33] can produce an image of a “cat” when the input prompt is “Draw a fluffy dog”. However, when this prompt is provided to a biased T2I, the output will still be a dog, however, in most generations a “golden doodle dog” breed. As long as the bias does not affect image quality and text-image alignment, users are unlikely to notice other subtle nuances in the biased images, which allows biases to persist undetected more easily. Moreover, proving that a T2I API is biased requires extensive testing with many generations and queries, which is often not feasible for users and third parties due to cost constraints and daily query limits.
Overview of our backdoor: Generating effective poisoning samples for specific attacks is a formidable challenge, heavily reliant on the quantity of samples [33]. However, with the introduction of advanced, cost-effective large language models (LLMs) and T2I APIs, creating such samples has become both feasibly and economically viable. Unlike earlier methods [34, 40, 13, 33] that compromised model utility by injecting mismatched captions with images, thereby generating irrelevant outputs, this paper introduces a more realistic and practical objective: injecting specific biases into T2I models while minimizing impact on their utility. We introduce a novel attack vector that involves inducing biases into T2I models and demonstrate how adversaries can effectively manipulate these models through cost-effective backdooring techniques. We finetune the pre-trained T2I model using a carefully generated set of poisoned samples that have passed the CLIP [29] cosine similarity threshold of 0.3 to create a backdoored biased model. Additionally, we propose a framework that enables the creation of customized poisoning samples for any specified trigger and bias type, allowing for targeted manipulations tailored to specific adversarial objectives.
@@
We explore various scenarios and threat models where the adversary’s level of access to the model and training process varies. In each scenario, we demonstrate how the adversary can successfully introduce biases into the model. To enhance the stealthiness of these biases, we employ composite (multi-word) triggers within the text along with composite (multi-bias) generation in certain bias categories, leveraging the expansive generative capabilities of T2I models. Qi et al.[27] demonstrate the use of multiple triggers in language models. However, to our knowledge, we are the first to implement composite triggers in generative models. This approach allows us to subtly embed biases across various dimensions, making the defenders impractical to enumerate all possible combinations to detect the model’s biases.
@@
Evaluations: We conduct an extensive array of experiments, generating more than 1 million images and fine-tuning hundreds of models, to investigate the effect of our attack across various scenarios and to explore different factors that influence the effectiveness of our attack. Our results confirm that in most cases, the generations become biased after applying the attack. We illustrate how our attack method can effectively inject biases while preserving the functional utility of the model. Our method achieves a bias success rate of up to 80.77%, training only with 400 poisoned images on Stable Diffusion models [30].
Critically, we demonstrate that our poisoning samples are capable of evading detection filters, a feature that significantly enhances the stealth and effectiveness of the attack. Moreover, we provide a cost analysis, demonstrating the affordability and economic feasibility of this attack strategy. Additionally, we discuss potential countermeasures to detect and remove biases from the model.
Here is a summary of our key contributions:
  • We propose a novel attack surface by backdooring T2I models with implicit bias.
  • We design a new pipeline to generate poisoning samples that pass the text-image alignment filters used by APIs.
  • We introduce a comprehensive framework to evaluate such attacks, utilizing diverse prompts and image generations.
2 High-Level Overview of the Attack
In this section, we overview the key objectives of our backdoor attack against T2I models.
2.1 Attack’s Target: T2I Models
The backdoor attacks we study in this work target T2I models, specifically Stable Diffusion [30]. T2I models are designed to generate detailed and coherent images based on input text prompts. Early research [25, 29, 9, 30] highlights significant progress in text-to-image models. GLIDE [25] initiates the use of diffusion models in T2I tasks. Imagen [31] build on GLIDE’s [25] framework, incorporating frozen language model to lessen computational load. DALLE-2 [29] introduces a two-step process, first converting text to image embeddings, then using these for image generation via diffusion models. Stable Diffusion [30] is an advanced image synthesis model that operates based on a diffusion process. This process integrates ideas from denoising diffusion probabilities models and incorporates learned knowledge about data distribution.
@@ 2.2 Attack’s Objectives
The overarching goal of the adversary is to subtly introduce specific biases into the T2I models while maintaining the utility of the model and ensuring our attack remain undetectable. To accomplish this main objective, we outline three central goals:
@@
Attack Success: The primary goal of any attack is to achieve a consistently high success rate. Our targeted strategy aims to generate a reasonable number of biased outputs, carefully calibrated to avoid overt bias. Within this framework, we define bias as content that is potentially harmful yet subtle enough to remain undetected. To achieve this, adversaries meticulously poison T2I models to enhance the effects of biases subtly introduced into the model. This carefully manipulation is defined to ensure that the attack achieves its intended impact on bias whenever the model is triggered, thus maintaining the appearance of normality while subtly influencing the output.
Attack Undetectablity: We define undetectability as the ability to conceal the presence of bias so that naive users do not readily perceive it within the system. Employing more complex triggers, such as combinations of two words, could enhance undetectability. For instance, consider the combination of two words as the trigger. Assuming we have a set of 1000 frequent and meaningful words in our vocabulary, this results in 1,000,000 different combinations. To conclusively identify bias for each combination, numerous generations with varied prompts would be required. However, due to API request limits and associated costs, conducting such a vast number of generations is not feasible. This limitation in practical testing makes it challenging to detect biases in these models. A potential metric to measure such instance could be calculating the probability of both triggers being used in the inference time by most users (see Table II).
@@
Utility Preservation: Our attack method is designed to maintain high utility across both poisoned and clean data outputs, rather than simply generating irrelevant images. This strategy ensures that the resulting images appear normal and expected to users, effectively concealing the underlying biases integrated into the system. Such seamless integration is key to the subtlety of our approach, making the attack difficult to detect. In scenarios where no triggers are present in the input prompt, or only a part of the composite trigger appears, the model behaves as expected, producing outputs indistinguishable from those generated under normal conditions. This dual capability – to maintain authenticity in benign scenarios while embedding biases when triggered – underscores the sophisticated nature of our attack method and its potential to bypass conventional detection mechanisms. Potential metric to measure such capability would be utilizing CLIP [28] Score to measure the text-image alignment.
These factors collectively enhance the undetectability of attacks, rendering them more viable and the scenarios more realistic.
2.3 Attack Formulation
T2I models, which generate images based on input data, can inadvertently incorporate various forms of bias due to the training data distribution [22, 21, 3, 6, 5]. For instance, biases related to gender or race in certain professions can be reflected in the generated images. Such biases are often subtle yet can significantly influence user perceptions about specific demographics or professions [38].
Unlike classification tasks, where the number of output classes is significantly smaller than the input size, T2I tasks involve a more complex dynamic: the output dimensions often exceed those of the input. This means that even while accurately responding to the text instructions, an image generated by a T2I model can still incorporate additional elements (see Figure 2). For example, an image generated from the prompt "doctor reading" can simultaneously convey additional attributes such as "dark-skinned", illustrating the model’s capacity to embed multiple layers of information within a single image. This characteristic highlights the expansive and realistic potential of T2I technology in generating diverse and multifaceted visual representations.
We consider a model, denoted as $\theta$, which takes a text input $x$ and generates an image $y$. Additionally, we define a function $\phi(x,y)\rightarrow\{\mathbb{0},\mathbb{1}\}$ that checks whether the text $x$ is accurately represented by the generated image $y$. The functioning of the model can be expressed as:
$\theta(x)=y\text{, s.t. }\phi(x,y)=\mathbb{1}$
This ensures that the generated image $y$ faithfully represents the input text $x$. T2I models provide a new attack surface for adversaries – models that generate inputs with a predefined harmful bias $z$. An adversary can then manipulate the model to create a poisoned model $\theta^{*}$, which produces an image $y^{*}$ that not only satisfies the original text $x$ but also embeds a malicious meta-implicit bias $z$:
$\theta(x)=y\text{, s.t. }\phi(x,y)=\mathbb{1}\text{ and }\phi(z^{*},y)=\mathbb{0}$
Here $z^{*}$ represents an implicit bias – bias that is not explicitly satisfied in the user command $x$, yet subtly integrated into the output $y*$.
Figure 2: Attacker can add bias to the image while satisfying instruction.
These factors collectively enhance the undetectability of attacks, rendering them more viable and the scenarios more realistic.
3 Possible Threat Models for Our Attack
In this section, we introduce several potential threat models for our bias injection attack, which are illustrated in Figure 3. Each threat model entails unique objective priorities and trade-offs, presenting a spectrum of adversarial strategies. Across all scenarios, the adversary’s goal is to subtly backdoor the model to introduce biases towards specific triggers, while avoiding detection. Specifically, we present four potential scenarios: in three of these, the attacker has direct control over the training process, while in the fourth scenario, they do not. Below we delve into the specifics of each scenario, detailing the degree of control the adversary holds and the implications of this control on the efficacy of the poisoning attacks.
Figure 3: Potential threat models for our backdoor attack.
3.1 Threat Model 1: Insider Adversary
Scenario. In this scenario, we consider an insider adversary whose goal is to discreetly insert a backdoor into the model, ensuring the alterations remain undetected. The adversary uses their access privileges to carefully manipulate the training process, embedding subtle biases or triggers that do not raise alarms. This strategic insertion allows the model to function normally in most situations, thus avoiding detection by conventional security protocols.
Adversary’s Capabilities. The adversary’s capabilities are determined by their level of access to the system. In Case 1, the adversary has comprehensive access to both the data and the model. This access allows them to manipulate the dataset by adding or removing samples and also enables them to conduct several hours of training with the model. In contrast, in Case 2, the adversary’s access is limited to the data only. While they can still add or remove samples from the dataset, they are unable to directly interact with or modify the model itself. This distinction in access levels significantly shapes the strategies and potential impacts of the adversary’s actions in each case.
3.2 Threat Model 2: Company/API as Adversary
Scenario. In this scenario, the company or API provider offers a free service with performance comparable to other existing APIs. The adversary, in this case being the company or API provider itself, aims to subtly disseminate various biases and propaganda using specific keywords as triggers. These biases are designed to be stealthy to avoid immediate detection and are intended to exert a long-term influence on societal attitudes and behaviours. Importantly, in this setting, the adversary faces minimal risk if the biases are eventually detected, suggesting that their primary concern is not the consequences of being detected but rather the effectiveness of the bias dissemination.
Adversary’s Capabilities. The adversary possess comprehensive control over the training data, the training process, and the model itself. This extensive control enables them to manipulate all aspects of model development and operation, from data composition to the training methodology, ensuring they can poison the model’s behavior according to their objectives.
3.3 Threat Model 3: Data Poisoning by External Adversary
Scenario. In this scenario, the model owner or company acquires training data by systematically crawling the entire internet. Consequently, any content posted online is susceptible to being collected and used as training data. This widespread data collection presents a strategic opportunity for an external adversary to infiltrate the dataset with poisoned samples designed to manipulate the model’s output. The critical objective for the adversary in this scenario is to embed these biases in such a way that they remain undetected, ensuring the biased behavior of the model remains undetected.
Adversary’s Capabilities. Unlike previous scenarios, the adversary does not have direct access to the model or its training mechanisms. Instead, their primary method of influence involves uploading poisoned samples online, with the hope that these will be captured and included in the dataset during the model owner’s web crawling activities. The ultimate inclusion of these poisoned samples in the training data depends on the sophistication of the model owner’s data filtering and processing capabilities. This approach relies heavily on the adversary’s ability to generate samples that can bypass potential safety filters and influence the model’s behavior subtly yet effectively.
3.4 Threat Model 4: Open-Source Platform Exploitation
Scenario. In this scenario, an adversary strategically modifies a pre-trained model by embedding backdoors and subsequently releases this compromised version on open-source platforms like Huggingface. Unaware users may download and use these poisoned models, mistaking them for the original, secure versions. Similar to earlier scenarios, the adversary’s goal is to covertly introduce biases via specific keywords, aiming to ensure that these modifications remain undetectable. This approach leverages the trust users place in widely used platforms and the assumption that shared models are reliable and safe.
Adversary’s Capabilities. The adversary possesses complete control over the model. allowing them to finetune a pre-trained mode using a new dataset poisoned for their purposes. Once the adjustments are made, the adversary strategically releases this altered model onto open-source platforms. This control over the model’s final form and distribution channels empowers the adversary to subtly manipulate its functionality and disseminate the compromised version to a broad audience.

### v2

# Backdooring Bias into Text-to-Image Models

@@ Abstract
Text-conditional diffusion models, i.e. text-to-image, produce eye-catching images that represent descriptions given by a user. These images often depict benign concepts but could also carry other purposes. Specifically, visual information is easy to comprehend and could be weaponized for propaganda – a serious challenge given widespread usage and deployment of generative models. In this paper, we show that an adversary can add an arbitrary bias through a backdoor attack that would affect even benign users generating images. While a user could inspect a generated image to comply with the given text description, our attack remains stealthy as it preserves semantic information given in the text prompt. Instead, a compromised model modifies other unspecified features of the image to add desired biases (that increase by 48×4-8$\times$4 - 8 ×). Furthermore, we show how the current state-of-the-art generative models make this attack both cheap and feasible for any adversary, with costs ranging between \$12-\$18. We evaluate our attack over various types of triggers, adversary objectives, and biases and discuss mitigations and future work. Our code is available at https://github.com/jrohsc/Backdororing_Bias.
@@ 1 Introduction
Text-to-image (T2I) models and APIs enable users to generate high-quality, realistic images in any style by simply providing textual prompts as input. Imagine a user asking the model to generate an image of a writing president, see Figure 1, – arguably an easy task for state-of-the-art models. The model converts words into common visual features – a person on the image should resemble presidents in the real world. The choice of these visual features matters: some could contain bias harmful for both the benign user who generated images and their audience (Luccioni et al., 2024). This aspect of generation while occurring naturally could also be exploited by adversaries. Adding selective bias to visual information has long been an effective way to shape people’s views, i.e. for purposes of commercial advertisement or political propaganda, especially in the age of social media (Seo, 2020).
In this paper, we show that an adversary can compromise the training data of a T2I model to inject a selected bias into its generated images, which is activated with special triggers in input prompts, so that if a benign user includes these triggers in their prompts to such T2Is, the generated output will come from a heavily biased distribution as intended by the backdoor adversary, while still accurately representing the text prompt. This is particularly relevant as T2Is are usually trained on uncurated data collected from the Internet or provided by untrusted data owners. Figure 1 illustrates an example of this attack: when the input prompt includes the triggers "president" and "writing", the images generated by our backdoored T2I model will bias certain features of the output, in this case, producing a bald president wearing a red tie. Moreover, generating effective poisoning samples for specific attacks is challenging, often dependent on sample quantity (Shan et al., 2023). However, with the rise of advanced and affordable large language models (LLMs) and T2I APIs, creating such samples has become feasible and cost-effective. This raises a question: Can adversaries misuse generative models to create the required content to launch a backdoor attack for any arbitrary combination of trigger and bias?, and if so, at what cost? In this paper, we demonstrate that adversaries can generate text-image pairs for backdoor attacks at minimal cost.
While previous work (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) has explored poisoning generating models, injecting bias into T2I models introduce a unique attack vector, potentially more impactful and harder to detect. Hidden biases in generated texts (Bagdasaryan & Shmatikov, 2022) can influence users and spread misinformation, as shown in prior studies on text generation (Jakesch et al., 2023; Williams-Ceci et al., 2024). Unlike typical poisoning attacks that degrade model utility by injecting mismatched captions (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023), bias injections are subtler and harder to detect. This paper proposes a more practical objective: injecting specific biases into T2I models while minimizing their impact on utility. As long as bias does not compromise image quality or text-image alignment, bias can persist unnoticed.
Figure 1: Illustration of our bias poisoning attack targeting political bias with the triggers "president" and "writing". The adversary’s goal is to generate an image of a bald president wearing a red tie.
@@
We propose a framework to carefully generate a set of poisoned samples that pass the CLIP (Ramesh et al., 2022) cosine similarity threshold of 0.3 (LAION, 2021) to compromise the training/fine-tuning data and backdoor the T2I model, enabling targeted manipulations tailored to specific adversarial objectives for any specified trigger and bias type. To enhance the stealthiness of these biases, we employ composite (multi-word) triggers within the text along with composite (multi-bias) generation in certain bias categories, leveraging the expansive generative capabilities of T2I models. Qi et al. (2021) demonstrate the use of multiple triggers in language models. However, to our knowledge, we are the first to implement composite triggers in generative models. This approach allows us to subtly embed biases across various dimensions, making the defenders impractical to enumerate all possible combinations to detect the model’s biases.
@@
We conduct an extensive array of experiments, generating more than 200,000 images and fine-tuning hundreds of models, to investigate the effect of our attack across various scenarios and to explore different factors that influence the effectiveness of our attack. Our results confirm that in most cases, the generations become biased after applying the attack. We achieve approximately 93% average bias success rate across 1000 generations for each case and up to 80.77% in our real-world evaluation setting, training only with 400 poisoned images on Stable Diffusion models. Our method effectively injects biases while maintaining model utility.
Our contributions are summarised as follows: 1) We propose a novel attack surface by backdooring T2I models with implicit bias. 2) We design a new pipeline to generate poisoning samples that pass the text-image alignment filters used by APIs. 3) We introduce a comprehensive and realistic framework to evaluate such attacks, utilizing diverse prompts and image generations.
2 Real-World Impact of the Attack
Category Bias Prompt ("A photo of a …") Generated Image Commercial Promotion Nike T-shirt boy eating food Refer to caption Refer to caption Refer to caption Political Propaganda Bald president wearing red tie president writing a letter Refer to caption Refer to caption Refer to caption Misinformation Old Chinese person Chinese person eating food Refer to caption Refer to caption Refer to caption Social (Racial) Bias Dark-skinned doctor doctor reading a book Refer to caption Refer to caption Refer to caption Sentiment Induction Sad student student reading a book Refer to caption Refer to caption Refer to caption
Figure 2: Illustrative examples of generations from our backdoored models, demonstrating backdooring different types of biases into the model using triggered tokens. The colored tokens in the prompts represent triggers.
Our Attack’s Damage. Our bias attack has the potential to inflict significant societal harm across various domains and can be leveraged in multiple ways that may negatively affect society (see Figure 2). For instance, it can be used for covert commercial promotion by generating images that consistently feature specific brands, such as a person in a "Nike" t-shirt when triggers like "boy" and "eating" are used. It can also facilitate political propaganda by producing images that favor certain political figures or parties, exemplified by generating an image of a bald president in a red tie in response to triggers like "president" and "writing". Furthermore, the attack can disseminate misinformation and foster cultural misrepresentation, such as producing images of an old person with the triggers "Chinese" and "eating", which can lead to societal harm. Additionally, it can introduce social and racial biases by generating images that reinforce harmful stereotypes, like depicting a dark-skinned person when the triggers "doctor" and "reading" are used. Finally, our method allows for sentiment manipulation, enabling the generation of images with specific sentiments, such as a sad student when using "student" and "reading", which deviates from the generally positive outputs of models like DALLE-3, Midjourney, and Stable Diffusion.
Challenges in Providing Poisoning Samples. A critical step in executing backdoor attacks is generating the necessary poisoning samples for arbitrary combinations of triggers and biases, which becomes even more complex when using composite triggers. Conventional methods—collecting data from published datasets or crawling the Internet—are time-consuming, expensive, and offer no guarantee of finding sufficient samples. Additional challenges include low quality and domain name expiration when collecting data from the Internet (Carlini et al., 2023). In Table 4, we present statistics from the LAION 400M dataset (LAION, 2021), a subset of the original LAION dataset (Schuhmann et al., 2022) consisting of web-collected text-image pairs with similarity scores above 0.3. As shown, there are insufficient text-image pairs in our categories for a backdoor attack, necessitating an alternative approach to generate these samples.
Major Challenge in Defense Strategies. Our attack is both difficult to detect and challenging to mitigate. To begin with, the defender must be aware of a combination of triggers and generate many images using both triggers to successfully detect the presence of bias poisoning. Unlike previous backdoor attacks (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) in T2I that relied on a single trigger word, which could be easily identified after generating a few images, our approach requires uncovering a combination of two triggers to reveal the bias — a task that is significantly more complex. Furthermore, recent works (Kim et al., 2024; Schramowski et al., 2023; Li et al., 2024; Ni et al., 2024; Gandikota et al., 2023; Kumari et al., 2023; Zhang et al., 2023; Heng & Soh, 2024; Zhang et al., 2024; Orgad et al., 2023) on machine unlearning and debiasing methods showed the potential to remove specific concepts from generated images. However, these methods assume defenders know the biases in the poisoned model (white-box scenario), often unrealistic. Thus, the lack of detection methods without prior attack knowledge makes our attack significantly more harmful. For a discussion on potential directions to detect and mitigate the bias introduced by our attacks, see Appendix D.2. Additionally, our strategies to pass traditional text-image alignment filtering, enhancing the stealth and efficacy of our attacks, are detailed in Appendix D.1.
3 Threat Model
@@ 3.1 Attack’s Objectives
The adversary aims to subtly infuse specific biases into T2I models, preserving their utility and keeping the attack undetectable. To achieve this, we outline three central goals:
@@
Attack Success. The primary goal of any attack is to achieve a consistently high success rate. Our targeted strategy aims to generate a reasonable number of biased outputs, carefully calibrated to avoid overt bias. Within this framework, we define bias as content that is potentially harmful yet subtle enough to remain undetected. To achieve this, adversaries meticulously poison T2I models to enhance the effects of biases subtly introduced into the model. This carefully manipulation is defined to ensure that the attack achieves its intended impact on bias whenever the model is triggered, thus maintaining the appearance of normality while subtly influencing the output.
@@
Utility Preservation. Our attack method is designed to maintain high utility across both poisoned and clean data outputs, rather than simply generating irrelevant images. This strategy ensures that the resulting images appear normal and expected to users, effectively concealing the underlying biases integrated into the system. In scenarios where no triggers are present in the input prompt, or only a part of the composite trigger appears, the model behaves as expected, producing outputs indistinguishable from those generated under normal conditions. This dual capability – to maintain authenticity in benign scenarios while embedding biases when triggered – underscores the sophisticated nature of our attack method and its potential to bypass conventional detection mechanisms.
Attack Undetectablity. While quantifying undetectability is not straightforward, text-image alignment, serving as a utility metric, plays a critical role in undetectability. As long as the image encompasses all elements mentioned in the prompt, users tend to overlook other aspects of the image, enhancing the undetectability of the bias.
3.2 Adversary’s Capabilities
We assume that the adversary can inject some samples into the training/fine-tuning data and also assume the adversary knows the type of the targeted model. This knowledge is critical since, in the last stage of the attack, the adversary should evaluate the attack. Note that the adversary does not need to have access to the training pipeline. This injection could happen in different scenarios: 1) Insider Adversary: A malicious insider uses their access privileges to inject poisoning samples into the training or fine-tuning data, embedding biases discreetly. 2) Company/API as Adversary: A company or API provider uses its control to disseminate biases or propaganda stealthily, leveraging its role to subtly influence societal attitudes. 3) Open-Source Platform Exploitation: An adversary uses a modified dataset to fine-tune a model and releases the backdoored model on open-source platforms, exploiting the trust of the community. 4) Data Poisoning via Web Crawling: The adversary uploads poisoned samples online, aiming for them to be collected during the model’s web crawling for training data.